OpenStudy (anonymous):

Use the coordinate plane illustrating quadrilaterals MNOP and EFGH to choose the correct statement. Rectangle MNOP is _ to rectangle EFGH. a) congruent b) similar c) no relation

OpenStudy (anonymous):

here

OpenStudy (anonymous):

is it c?

OpenStudy (saifoo.khan):

@amistre64 @satellite73

OpenStudy (anonymous):

I don't think that those are options ;)

jhonyy9 (jhonyy9):

i think b.

OpenStudy (anonymous):

@jhonyy9 How, the coordinate numbers don't look alike or related to each other at all?

OpenStudy (anonymous):

or not...

OpenStudy (amistre64):

congruent would be exact, barring rotation of course; so that ones out

OpenStudy (amistre64):

notice the the rotation is 45degrees; giving us a short side of sqrt(2) as compared with a short side of 3

OpenStudy (amistre64):

3n = sqrt(2); when n=sqrt(3) if the long sides are sclaed by n as well its similar; otherwise its neither

OpenStudy (amistre64):

when n=sqrt(2)/3 .... i was sure i typed that correctly :)

OpenStudy (anonymous):

oh, I thought it was C

OpenStudy (amistre64):

does\[5*\frac{\sqrt{2}}{3}=3\sqrt{2}\]

OpenStudy (anonymous):

Thank you for explaining!

OpenStudy (amistre64):

hope it helped :) but yes, I do believe its c